Find the probability that x is between 5 and 15.(round your answer to 4 decimal places)

X~N(8,2)
Answer #1

Answer:

Given,

Mean = 8

Standard deviation = 2

P(5 < X < 15) = P((5 - 8)/2 < (x-u)/s < (15 - 8)/2)

= P(-1.5 < z < 3.5)

= P(z < 3.5) - P(z < - 1.5)

= 0.9997674 - 0.0668072 [since from z table]

= 0.9329602

= 0.9330

Probability = 0.9330